OpenStudy (anonymous):

PLEASE HELP!!!!!!!!!!!!!!!!!!!!!!!!! MEDAL AWARDED Find the length of a segment that has endpoints H(2, -3) and J(12, 7). (1 point) A. 20 B. 10√2 C. 4√2 D. 40

jigglypuff314 (jigglypuff314):

use the distance formula which is: \[\sqrt{(x _{2}-x _{1})^{2}+(y _{2}-y _{1})^{2}}\] given the points (x1, y1) and (x2, y2) so plug in x1 = 2 , y1 = -3 , x2 = 12 , y2 = 7 and solve, which will give you your distance :)

OpenStudy (anonymous):

Okay one sec let me work it out

OpenStudy (anonymous):

(12 -2) ^2 + (12 -(-3) ^2 right?

jigglypuff314 (jigglypuff314):

(12-2) ^2 + ((7) -(-3)^2 almost :) and don't forget to square root it

OpenStudy (anonymous):

I got 14.14 Did i do it right?

jigglypuff314 (jigglypuff314):

that is correct but I'm thinking from the options you might want to keep it with the square root so you got √200 = √(100*2) = √100 * √2 = 10√2

OpenStudy (anonymous):

Yes so the answer would be B right?

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (anonymous):

Thanks so much i couldn't have done it without you!!!

jigglypuff314 (jigglypuff314):

your welcome :) glad I could help :D